OpenStudy (anonymous):

write the slope intercept equation for the line that passes through (8,12) and is perpendicular to -4x-7y=14

OpenStudy (anonymous):

Find the slope of the line -4x-7y=14 by solving for y y = -(4/7)x -2 The slope is the -4/7 perpendicular lines have slopes that are negative reciprocals, The slope for the new line will be 7/4 Then use this new slope and the point (8,12) to find "b", the y-intercept. If you use the point - slope form of an equation of a line we get: y-12 = 7/4(x-8) which becomes y = 7/4 x -2 If I did the algebra correctly. OK?